### v2

# Affine complements in the projective space with the trivial group of automorphisms

Affine complements in the projective space
with the trivial group of automorphisms
A.V.Pukhlikov
We construct a new class of affine complements ${\mathbb{P}}^{M}\setminus S$ with the trivial group of automorphisms, where $S\subset{\mathbb{P}}^{M}$ is a rational hypersurface, $M$ is odd and $M\geqslant 5$.
Bibliography: 19 items.
Key words: affine complement, birational map, exceptional divisor.
AMS classification: 14E05, 14E07
Introduction
0.1. Automorphisms of affine complements. Let $X$ be a non-singular projective rationally connected variety, $Y\subset X$ a prime ample divisor. There is a natural problem of computing the group of automorphisms $\mathop{\rm Aut}(X\setminus Y)$ of the affine variety $X\setminus Y$. This problem can be seen both as a problem of affine algebraic geometry and as a problem of birational geometry: the automorphisms of the complement $X\setminus Y$ are birational self-maps of the variety $X$, regular on the open set $X\setminus Y$ and transforming it into itself. Of course, the group of biregular automorphisms $(\mathop{\rm Aut}X)_{Y}$, transforming $Y$ into itself, is a subgroup of the group $\mathop{\rm Aut}(X\setminus Y)$, and we could say that the problem of computing the latter group is (practically) solved, if the equality
$\mathop{\rm Aut}(X\setminus Y)=(\mathop{\rm Aut}X)_{Y}$
holds, which is of course true if the variety $Y$ is not birationally ruled (see [1, Proposition 1]). Since in any case
$\mathop{\rm Aut}(X\setminus Y)\subset\mathop{\rm Bir}X,$
the problem of computing the group of automorphisms of an affine complement is non-trivial only if the group $\mathop{\rm Bir}X$ of birational self-maps of the variety $X$ is sufficiently large, in particular, infinite. By what was said, it makes sense to consider only pairs $X\supset Y$, such that $X$ has a large (even very large) group of birational self-maps, and the ample divisor $Y$ is a birationally ruled variety. In the present paper, as in [1], we consider the case when $X={\mathbb{P}}^{M}$ is the complex projective space and $Y=S$ is a rational hypersurface. In order to state the main result, let us describe the family of these hypersurfaces.
0.2. Rational hypersurfaces. Fix the integers $e\geqslant 2$ and $m\geqslant e+1$. Set $M=2e+1$. Consider the complex projective space ${\mathbb{P}}={\mathbb{P}}^{M}$ of dimension $M\geqslant 5$ with the homogeneous coordinates
$(x_{0}:\dots:x_{e}:y_{0}:\dots:y_{e})$
(the meaning of these non-standard notations will become clear below). By the symbol $q_{\alpha,\beta}(x_{*},y_{*})$ we denote a bi-homogeneous polynomial in these coordinates of degree $\alpha\in{\mathbb{Z}}_{+}$ in $x_{*}$ and $\beta\in{\mathbb{Z}}_{+}$ in $y_{*}$. Obviously, such polynomials form a linear subspace of the space of homogeneous polynomials of degree $\alpha+\beta$ on ${\mathbb{P}}^{M}$. The symbol ${\cal P}(m)$ stands for the linear space of polynomials of the form
$q_{m,m+1}(x_{*},y_{*})+q_{m+1,m}(x_{*},y_{*}).$
By the symbol ${\cal P}_{\rm sym}(m)$ we denote the linear subspace in ${\cal P}(m)$, consisting of polynomials, symmetric with respect to the replacement $x_{i}\leftrightarrow y_{i}$, $i=0,\dots,e$, that is,
$q_{m+1,m}(y_{*},x_{*})=q_{m,m+1}(x_{*},y_{*}).$
Finally, let
$P_{1}=\{y_{0}=\dots=y_{e}=0\},\quad P_{2}=\{x_{0}=\dots=x_{e}=0\}$
be subspaces in ${\mathbb{P}}^{M}$. Obviously, $(x_{0}:\dots:x_{e})$ and $(y_{0}:\dots:y_{e})$ form natural systems of homogeneous coordinates on $P_{1}\cong{\mathbb{P}}^{e}$ and $P_{2}\cong{\mathbb{P}}^{e}$. For $f\in{\cal P}(m)$ we have $P_{i}\subset S(f)=\{f=0\}$, $i=1,2$, and moreover, if both components $q_{m,m+1}$ and $q_{m+1,m}$ are not identically zero, then for $i=1,2$ the equality
$\mathop{\rm mult}\nolimits_{P_{i}}S(f)=m$
holds. Let $p_{i}\in P_{i}$ be points of general position and $[p_{1},p_{2}]\subset{\mathbb{P}}$ the line going through these points. The restriction of the polynomial $f$ onto that line gives the divisor
$\left(f|_{[p_{1},p_{2}]}\right)=mp_{1}+mp_{2}+R(p_{1},p_{2}),$
where $R(p_{1},p_{2})\in S(f)$ is some point. We get a map
$R\colon P_{1}\times P_{2}\dashrightarrow S(f),$
which is birational (the inverse map $R^{-1}\colon S(f)\dashrightarrow P_{1}\times P_{2}$ is $\pi_{2}\times\pi_{1}|_{S(f)}$, where $\pi_{i}\colon{\mathbb{P}}\dashrightarrow P_{j}$ is the linear projection from the subspace $P_{i}$, $\{i,j\}=\{1,2\}$). Therefore, the hypersurface $S(f)$ is rational.
0.3. Statement of the main result. Let $\sigma\colon{\mathbb{P}}^{+}\to{\mathbb{P}}$ be the blow up of the disjoint union $P_{1}\cup P_{2}$ and $\Delta_{i}=\sigma^{-1}(P_{i})$ the non-singular exceptional divisors,
$\Delta_{i}\cong P_{i}\times{\mathbb{P}}^{e}\cong{\mathbb{P}}^{e}\times{\mathbb{P}}^{e},$
where
$((x_{0}:\dots:x_{e}),(y_{0}:\dots:y_{e}))$
are bi-homogeneous coordinates on $\Delta_{1}$ and
$((y_{0}:\dots:y_{e}),(x_{0}:\dots:x_{e}))$
are bi-homogeneous coordinates on $\Delta_{2}$. Here the first group of coordinates corresponds to $P_{i}$ and the second one to the direct factor. Using the same symbols as bi-homogeneous coordinates on $\Delta_{i}$, we somewhat abuse our notations, however, the second group of coordinates corresponds naturally to precisely the group of coordinates on ${\mathbb{P}}$ that complements the coordinates on $P_{i}$, so that these notations cannot lead to any misunderstanding. In terms of these systems of bi-homogeneous coordinates the intersection $S^{+}\cap\Delta_{1}$ is given by the equation $q_{m+1,m}(x_{*},y_{*})=0$, and the intersection $S^{+}\cap\Delta_{2}$ by the equation $q_{m,m+1}(x_{*},y_{*})=0$ (where $S^{+}\subset{\mathbb{P}}^{+}$ is the strict transform of the hypersurface $S=S(f)$). It is easy to show (see Subsection 1.1), that for a general polynomial $f\in{\cal P}(m)$ (or ${\cal P}_{\rm sym}(m)$) the hypersurface $S^{+}$ and its intersections $S^{+}\cap\Delta_{i}$, $i=1,2$, are non-singular. Denote by the symbols ${\cal S}$ and ${\cal S}_{\rm sym}$ the linear systems
$\{S(f)\,|\,f\in{\cal P}(m)\}\quad\mbox{and}\quad\{S(f)\,|\,f\in{\cal P}_{\rm sym}(m)\}$
of divisors on ${\mathbb{P}}$, respectively.
Set
${\cal X}\subset{\mathbb{P}}(H^{0}({\mathbb{P}}^{e},{\cal O}_{{\mathbb{P}}^{e}}(m)))$
to be the open $\mathop{\rm Aut}{\mathbb{P}}^{e}$-invariant set of non-singular hypersurfaces of degree $m$ in ${\mathbb{P}}^{e}$. Since the group of automorphisms of a smooth hypersurface of degree $m$ in ${\mathbb{P}}^{e}$ is finite [2], the action of the $(e+1)^{2}-1=(e^{2}+2e)$-dimensional group $\mathop{\rm Aut}{\mathbb{P}}^{e}$ fibres ${\cal X}$ into $(e^{2}+2e)$-dimensional orbits.
Let $\sigma_{S}=\sigma|_{S^{+}}\colon S^{+}\to S$ be the restriction of the blow up $\sigma$ onto $S^{+}$, that is, the blow up of the singular set $P_{1}\cup P_{2}$ on $S$, so that
$\sigma^{-1}_{S}(P_{i})=S^{+}\cap\Delta_{i}$
are the exceptional divisors of that blow up. Assuming that they are non-singular, we get that for a point $p\in P_{i}$ of general position $\sigma^{-1}_{S}(p)=\sigma^{-1}_{S}(p)\cap S^{+}$ is a non-singular hypersurface of degree $m$ in ${\mathbb{P}}^{e}$, that is, a point of the open set ${\cal X}$, which defines the maps
$\xi_{i}\colon P_{i}\dashrightarrow{\cal X},\quad i=1,2.$
The closure of the image of $P_{i}$ in ${\cal X}$ we denote by the symbol $\xi_{i}(P_{i})$. Since $\mathop{\rm dim}P_{i}=e$ and $\mathop{\rm dim}{\cal X}={m+e\choose e}-1\geqslant{2e+1\choose e}-1$, the following condition of general position for the polynomial $f$ (or the hypersurface $S$) makes sense:
()(*)( ∗ ) for some point $p\in P_{i}$, for which the hypersurface $\sigma^{-1}_{S}(p)\subset{\mathbb{P}}^{e}$ is non-singular, the image $\xi_{i}(P_{i})$ intersects the $\mathop{\rm Aut}{\mathbb{P}}^{e}$-orbit of the point $\xi_{i}(p)$ at this point only, $i=1,2$.
Obviously, if the hypersurface $S$ satisfies the condition ()(*)( ∗ ), then this condition holds for every point of general position $p\in P_{i}$.
Now we can state the main result of the present paper.
Theorem 0.1. Assume that the strict transform $S^{+}$ of the hypersurface $S\in{\cal S}$ is non-singular and moreover, the exceptional divisors $\sigma^{-1}_{S}(P_{i})$ are non-singular, $i=1,2$.
(i) If $S^{\prime}\subset{\mathbb{P}}$ is an irreducible hypersurface of degree $12m+$ 1, such that there is an isomorphism
$\chi\colon{\mathbb{P}}\backslash S\to{\mathbb{P}}\backslash S^{\prime}$
[…]
(iii) If $S\in{\cal S}_{\rm sym}$ satisfies the condition ()(*)( ∗ ), then
$\mathop{\rm Aut}({\mathbb{P}}\backslash S)=\langle\tau\rangle\cong{\mathbb{Z}}/2{\mathbb{Z}},$
where the projective involution $\tau$ is defined by the equalities
$\tau^{*}x_{j}=y_{j},\quad\tau^{*}y_{j}=x_{j},\quad j=0,1,\dots,e.$
0.4. The structure of the paper. The paper is organized in the following way. In §1 we prepare to the proof of the main theorem. In Subsection 1.1 we consider the singularities of the hypersurface $S$, justifying the assumptions of the theorem. In Subsection 1.2 the proof starts: in the assumption that the isomorphism of the affine varieties ${\mathbb{P}}\setminus S$ and ${\mathbb{P}}\setminus S^{\prime}$ is the restriction of a birational, but not a biregular, automorphism of the projective space, we resolve the singularities of the latter map and define the key objects and parameters, characterizing this map, and prove an analog of [1, Proposition 2] that gives a relation between these parameters. Besides, we note that the result of [1] can be made somewhat stronger (no changes in the proof are needed).
In §2 we prove the main theorem. It is done in a few steps. To begin with, we consider the sequence of blow ups, resolving the exceptional divisor $T$, corresponding to the hypersurface $S^{\prime}$ (Subsection 2.1). After that, we exclude one by one all options for the centre of the divisor $T$ on ${\mathbb{P}}$: when it is not contained in the singular set of the hypersurface $S$ and when it is contained in one of the subspaces $P_{1}$ or $P_{2}$ (Subsections 2.2–2.4). This proves the claim (i) of the main theorem. In Subsection 2.5 we prove the claims (ii) and (iii), which completes the proof of Theorem 0.1.
0.5. Remarks. Studying the groups of automorphisms of affine algebraic varieties is a classical and established topic in algebraic geometry. As usual, in the dimension 2 quite a lot is known: [3, 4] are dealing with groups of automorphisms of affine surfaces (as far as the author knows, V.A.Iskovskikh also had a proof of the theorem on the group of automorphisms of the affine plane ${\mathbb{A}}^{2}$, which, it seems, has never been published). The two-dimensional case is still a field of active research, see, for instance, [5, 6, 7, 8].
In the dimension 3 and higher we know a lot less. There are just a few papers with complete results on the groups of automorphisms of affine varieties (in the cases when the problem is not trivial), see [1, 9]. Such classical problems as computing the group of automorphisms of the complement in ${\mathbb{P}}^{3}$ to a non-singular cubic surface (see [10]) or the group of automorphisms of the affine space ${\mathbb{A}}^{3}$ remain open.
The problems of describing the groups of automorphisms of affine varieties are on the border with such areas of algebraic geometry as describing the groups of birational self-maps and their subgroups and the theory of birational rigidity. A considerable recent progress in these areas (see, for instance, [11, 12, 13, 14, 15, 16, 17]), gives a basis for some optimism in respect of the groups of automorphisms of affine varieties, too.
1 Preliminary constructions
In this section we consider the singularities of a general hypersurface $S$ in more detail (Subsection 1.1) and start the proof of the main theorem (Subsection 1.2).
1.1. Singularities of the hypersurface $S$. Before starting to prove Theorem 0.1, let us consider in more detail the linear systems ${\cal S}$ and ${\cal S}_{\rm sym}$, defined in Subsection 0.3. For their strict transforms on ${\mathbb{P}}^{+}$ we have the equalities
${\cal S}^{+}=\{\sigma^{*}S(f)-m\Delta_{1}-m\Delta_{2}\,|\,f\in{\cal P}(m)\}$
and
${\cal S}^{+}_{\rm sym}=\{\sigma^{*}S(f)-m\Delta_{1}-m\Delta_{2}\,|\,f\in{\cal P}_{\rm sym}(m)\},$
respectively.
Proposition 1.1. The following equality holds: $\mathop{\rm Bs}{\cal S}^{+}=\mathop{\rm Bs}{\cal S}^{+}_{\rm sym}=\emptyset$.
Proof. Since the polynomials $q_{m+1,m}$ and $q_{m,m+1}$ are arbitrary, these linear systems have no base points on $\Delta_{1}\cup\Delta_{2}$, whereas for any point $p\in{\mathbb{P}}\backslash(P_{1}\cup P_{2})$ and a general polynomial $f$ we have $f(p)\neq 0$. Q.E.D. for the proposition.
It follows that
$\mathop{\rm Bs}{\cal S}=\mathop{\rm Bs}{\cal S}_{\rm sym}=P_{1}\cup P_{2}$
and for a general divisor $S\in{\cal S}$ (or ${\cal S}_{\rm sym}$) its strict transform $S^{+}\subset{\mathbb{P}}^{+}$ is non-singular. It is also obvious that for a general divisor $S\in{\cal S}$ (or ${\cal S}_{\rm sym}$) the intersections $S^{+}\cap\Delta_{i}$ are non-singular divisors of bi-degree $(m+1,m)$ on $\Delta_{i}\cong{\mathbb{P}}^{e}\times{\mathbb{P}}^{e}$. Since $m\geqslant e+1$, they are not rationally connected (and even uniruled) varieties.
1.2. The birational self-map $\chi_{\mathbb{P}}$. Let us start the proof of Theorem 0.1. Let
$\chi\colon{\mathbb{P}}\backslash S\to{\mathbb{P}}\backslash S^{\prime}$
be a non-trivial isomorphism of affine varieties, that is, the corresponding birational self-map $\chi_{\mathbb{P}}\in\mathop{\rm Bir}{\mathbb{P}}$ is not a biregular automorphism of the projective space. Consider the composition
$\chi_{\mathbb{P}}\circ\sigma\colon{\mathbb{P}}^{+}\dashrightarrow{\mathbb{P}}.$
Let $\varphi\colon\widetilde{\mathbb{P}}\to{\mathbb{P}}^{+}$ be a resolution of singularities of that map (we will see below that $\chi_{\mathbb{P}}\circ\sigma$ can not be regular, but for the moment we assume that if $\chi_{\mathbb{P}}\circ\sigma$ is regular, then $\varphi$ is the identity map). Let
$T=(\chi_{\mathbb{P}}\circ\sigma\circ\varphi)^{-1}_{*}S^{\prime}\subset\widetilde{\mathbb{P}}$
be the strict transform of the hypersurface $S^{\prime}$. Since $\chi_{\mathbb{P}}\not\in\mathop{\rm Aut}{\mathbb{P}}$, we see that the image $(\sigma\circ\varphi)(T)\subset{\mathbb{P}}$ is not the hypersurface $S$, so that $T$ is a $(\sigma\circ\varphi)$-exceptional divisor. Set
$B_{\mathbb{P}}=(\sigma\circ\varphi)(T)\subset{\mathbb{P}}.$
This is a subvariety of codimension $\geqslant 2$. Set also $B=\varphi(T)\subset{\mathbb{P}}^{+}$. It is possible that $B=\Delta_{1}$ or $\Delta_{2}$, but if this is not the case, then $B\subset{\mathbb{P}}^{+}$ is a subvariety of codimension $\geqslant 2$. Obviously, $B_{\mathbb{P}}\subset S$ and
$B\subset S^{+}\cup\Delta_{1}\cup\Delta_{2},$
where $\sigma(B)=B_{\mathbb{P}}$. Let
$a=a(T,{\mathbb{P}})\geqslant 1\quad\mbox{and}\quad a_{+}=a(T,{\mathbb{P}}^{+})\geqslant 0$
be the discrepancies and
$b=\mathop{\rm ord}\nolimits_{T}\varphi^{*}\sigma^{*}S\quad\mbox{and}\quad b_{+}=\mathop{\rm ord}\nolimits_{T}\varphi^{*}S^{+}.$
If $T=\Delta_{1}$ or $\Delta_{2}$, then $a_{+}=b_{+}=0$, otherwise $a_{+}\geqslant 1$. If $B\not\subset\Delta_{1}\cup\Delta_{2}$, then $a=a_{+}$ and $b=b_{+}$; in any case $b\geqslant 1$. If $B\subset\Delta_{1}\cup\Delta_{2}$, then since $\Delta_{1}\cap\Delta_{2}=\emptyset$, there is precisely one value $i\in\{1,2\}$, such that $B\subset\Delta_{i}$, and in that case set
$\varepsilon=\mathop{\rm ord}\nolimits_{T}\varphi^{*}\Delta_{i}.$
If $B\not\subset\Delta_{1}\cup\Delta_{2}$, then set $\varepsilon=0$.
We get the obvious equalities
$a=a_{+}+\varepsilon e,\quad b=b_{+}+\varepsilon m.$
Proposition 1.2. The following equality holds:
$(M+1)b=a(2m+1)+(M+1).$
Proof repeats the proof of the second equality (3.1) in [1] almost word for word. Namely, let $\widetilde{S}\subset\widetilde{\mathbb{P}}$ be the strict transform of the hypersurface $S$ and $\widetilde{K}=K_{\widetilde{\mathbb{P}}}$ the canonical class. We get the equalities
$\widetilde{K}=-(M+1)H+aT+(\dots),$
$\widetilde{S}\sim(2m+1)H-bT+(\dots),$
where for simplicity we write $H$ instead of $\varphi^{*}\sigma^{*}H$ and in the brackets we have integral linear combinations of $(\sigma\circ\varphi)$-exceptional divisors that are different from $T$. Their images with respect to the morphism $\chi_{\mathbb{P}}\circ\sigma\circ\varphi$ are of codimension $\geqslant 2$ in ${\mathbb{P}}$ and for that reason a general line $L\subset{\mathbb{P}}$ does not meet them. Furthermore, $\chi_{\mathbb{P}}\circ\sigma\circ\varphi(\widetilde{S})$ also is of codimension $\geqslant 2$ in ${\mathbb{P}}$. Therefore, for the strict transform $\widetilde{L}\subset\widetilde{\mathbb{P}}$ the following equalities hold:
$(\widetilde{L}\cdot\widetilde{S})=0,\quad(\widetilde{L}\cdot\widetilde{K})=-(M+1),$
and $(\widetilde{L}\cdot T)=(L\cdot S^{\prime})=2m+1$. Set $d=(\widetilde{L}\cdot H)$, it is the degree of the curve $\sigma\circ\varphi(\widetilde{L})\subset{\mathbb{P}}$ in the usual sense. As a result we get
$(2m+1)d-b(2m+1)=0,$
so that $d=b$, and
$-(M+1)d+a(2m+1)=-(M+1),$
which completes the proof of the proposition.
Now let us show that $T\neq\widetilde{\Delta_{1}},\widetilde{\Delta_{2}}$, so that the map $\chi_{\mathbb{P}}\circ\sigma$ can not be regular. Assume that $T=\widetilde{\Delta_{1}}$. In the notations of the proof of Proposition 1.2 we get $d=b=m$, so that the curve $\sigma\circ\varphi(\widetilde{L})$ is a rational curve of degree $m$ in ${\mathbb{P}}$, the strict transform of which $\varphi(\widetilde{L})\subset{\mathbb{P}}^{+}$ meets $\Delta_{1}$ at $(2m+1)$ points of general position. This is impossible: take a general hyperplane $\Xi\supset P_{1}$, then $(\Xi\cdot[\sigma\circ\varphi(\widetilde{L})])=m$ and $(\Xi^{+}\cdot(\varphi(\widetilde{L}))=-m-1$, since $\sigma^{*}\Xi=\Xi^{+}+\Delta_{1}$, where $\Xi^{+}\subset{\mathbb{P}}^{+}$ is the strict transform.
Remark 1.1. In [1] for the hypersurface $S$ a hypersurface of degree $\mathop{\rm deg}S\geqslant M+1$ with a unique singular point $o\in S$ of multiplicity $\mathop{\rm deg}S-1$, resolved by one blow up, was considered. The notations of [1] are close to the notations of the present paper, and where they are identical, they can be used in both contexts. In [1] in the proof of regularity of the map $\chi_{\mathbb{P}}$ we missed the case, similar to the one that we have just considered, when the hypersurface $S^{\prime}$ is transformed to the the exceptional divisor of the blow up of the point $o$ on ${\mathbb{P}}={\mathbb{P}}^{M}$. That case is easy to exclude by the arguments similar to our arguments above: the strict transform of a general line with respect to $\chi_{\mathbb{P}}$ is a rational curve of degree $\mathop{\rm mult}_{o}S=\mathop{\rm deg}S-1$ with the multiplicity $\mathop{\rm deg}S^{\prime}=\mathop{\rm deg}S$ at the point $o$, which is impossible. Note that the claim of Theorem 2 in [1] holds for any irreducible hypersurface $S^{\prime}\subset{\mathbb{P}}$ of degree $\mathop{\rm deg}S$, since the proof does not use any other information about $S^{\prime}$. The assumptions about that hypersurface made in [1, Subsection 1.1] are unnecessary.
2 Proof of the main result
In this section we prove Theorem 0.1. To begin with, we consider the sequence of blow ups, resolving the exceptional divisor $T$, and certain invariants, related to this sequence (Subsection 2.1). Then we exclude one by one the non-singular case, when $\sigma(B)\not\subset\mathop{\rm Sing}S$ (Subsection 2.2) and the singular case, when $\sigma(B)\subset P_{1}\cup P_{2}$ (Subsections 2.3 and 2.4), which completes the proof of the part (i) of Theorem 0.1. In Subsection 2.5 we prove the claims (ii) and (iii).
2.1. The resolution of the exceptional divisor $T$. As we saw in §1, $T$ is a divisor over ${\mathbb{P}}^{+}$, since $T\neq\Delta_{1},\Delta_{2},S^{+}$. Its centre on ${\mathbb{P}}^{+}$ is an irreducible subvariety $\varphi(T)\subset S^{+}\cup\Delta_{1}\cup\Delta_{2}$ of codimension $\geqslant 2$. Let
$\varphi_{i,i-1}\colon X_{i}\to X_{i-1},$
$i=1,\dots,N$, be the resolution of the divisor $T$ in the sense of [18, Chapter 2], that is, $X_{0}={\mathbb{P}}^{+}$, the birational morphism $\varphi_{i,i-1}$ is the blow up of the centre of $T$ on $X_{i-1}$, which we denote by the symbol $B_{i-1}$, where $\mathop{\rm codim}(B_{i-1}\subset X_{i-1})\geqslant 2$ for $i=1,\dots,N$, with the exceptional divisor $E_{i}=\varphi^{-1}_{i,i-1}(B_{i-1})$ and $E_{N}\subset X_{N}$ is a realization of the discrete valuation, corresponding to $T$, that is, the birational map
$\varphi^{-1}_{N,N-1}\circ\dots\circ\varphi^{-1}_{1,0}\circ\varphi\colon\widetilde{{\mathbb{P}}}\dashrightarrow X_{N}$
is regular at the general point of $T$ and transforms $T$ into $E_{N}$. We use the standard notations: for $j>i$ the composition
$\varphi_{i+1,i}\circ\dots\circ\varphi_{j,j-1}\colon X_{j}\to X_{i}$
is denoted by the symbol $\varphi_{j,i}$, and the strict transform of a subvariety $R\subset X_{i}$ on $X_{j}$, if it is well defined, is denoted by the symbol $R^{j}$. Note that the blown up subvarieties $B_{i-1}$, generally speaking, can be singular, however
$B_{i-1}\not\subset\mathop{\rm Sing}X_{i-1}$
for all $i=1,\dots,N$, and over a general point of $B_{i-1}$ the variety $X_{i}$ and the exceptional divisor $E_{i}$ are non-singular. By construction, for $j>i$ we have
$\varphi_{j-1,i-1}(B_{j-1})=B_{i-1}.$
On the set of exceptional divisors $E_{1},\dots,E_{N}$ there is a natural structure of an oriented graph: an oriented edge (an arrow) goes from $E_{i}$ to $E_{j}$ (which is written down as $i\to j$), if and only if $i>j$ and $B_{i-1}\subset E^{i-1}_{j}$. If $E_{i}$ and $E_{j}$ are not joined by an oriented edge, we write $i\nrightarrow j$.
The number of paths from the vertex $E_{i}$ to $E_{j}$ for $i\neq j$ in the oriented graph described above we denote by the symbol $p_{ij}$. Obviously, for $i\neq j$ we have $p_{ij}\geqslant 1$ if and only if $i>j$, and for $i<j$ we have $p_{ij}=0$. For convenience we set $p_{ii}=1$. Obviously, for $i>j$ we have the equality
$p_{ij}=\mathop{\rm ord}\nolimits_{E_{i}}\varphi^{*}_{i,j}E_{j}.$
(It justifies the convention $p_{ii}=1$.) Besides, in order to simplify the notations, we set $p_{i}=p_{Ni}$. (See [18, Chapter 2] or [19, Chapter 2] for the details.) Finally, for $i=1,\dots,N$ set
$\delta_{i}=\mathop{\rm codim}(B_{i-1}\subset X_{i-1})-1.$
Now we get the formula for the discrepancy
$a_{+}=\sum^{N}_{i=1}p_{i}\delta_{i}.$
Set also
[…]
where for simplicity of notations we write $S^{i-1}$ instead of $(S^{+})^{i-1}$; if the set in the right hand side is empty, we set $k=0$. Since the hypersurface $S^{+}$ is non-singular, for $i=1,\dots,k$ the hypersurface $S^{i-1}$ is non-singular at the general point of $B_{i-1}$, so that we get the formula
$b_{+}=\sum^{k}_{i=1}p_{i},$
which remains true for $k=0$, since in that case $b_{+}=0$.
2.2. Exclusion of the non-singular case. We say that the non-singular case takes place, if
$B_{\mathbb{P}}=\sigma(B)\not\subset\mathop{\rm Sing}S=P_{1}\cup P_{2},$
that is, $B\not\subset\Delta_{1}\cup\Delta_{2}$.
Proposition 2.1. The non-singular case does not realize: $B\subset\Delta_{i}$ for some $i\in\{1,2\}$.
Proof. Assume the converse: $B_{\mathbb{P}}\not\subset\mathop{\rm Sing}S$. Then $a=a_{+}$ and $b=b_{+}$, so that from Proposition 1.2, taking into account the explicit formulas above, we get the equality
$(M+1)\sum^{k}_{i=1}p_{i}=(2m+1)\sum^{N}_{i=1}p_{i}\delta_{i}+(M+1).$
Since $\delta_{i}\geqslant 1$, $22m+1\geqslant 2e+3\geqslant M+$ 2 and $k\leqslant N$, this equality is impossible: the right hand side is strictly higher than the left hand one. This contradiction proves the proposition. Q.E.D.
Therefore, $B\subset\Delta_{1}\cup\Delta_{2}$. Since $\Delta_{1}\cap\Delta_{2}=\emptyset$, the subvariety $B$ is contained in precisely one of the divisors $\Delta_{1},\Delta_{2}$, which we now denote by the symbol $\Delta$. Set
[…]
By construction, $l\geqslant 1$. Since the subvariety $\Delta$ is non-singular, we get the formula
$\varepsilon=\mathop{\rm ord}\nolimits_{T}\varphi^{*}\Delta=\sum^{l}_{i=1}p_{i},$
so that Proposition 1.2 gives the equality
$(M+1)\left[\sum^{k}_{i=1}p_{i}+m\sum^{l}_{i=1}p_{i}\right]=(2m+1)\left[\sum^{N}_{i=1}p_{i}\delta_{i}+e\sum^{l}_{i=1}p_{i}\right]+(M+1).$ (1)
2.3. Exclusion of the singular case for $l\leqslant k$. The following fact is true (here we do not assume that $l\leqslant k$).
Proposition 2.2. Assume that $k\geqslant 1$. Then for any $i\leqslant\mathop{\rm min}\{l,k\}$, $i\geqslant 1$, the inequality $\delta_{i}\geqslant 2$ holds.
Proof. By construction, the subvariety $B_{i-1}$ is contained in $E_{i-1}$, and by the inequalities $i\leqslant l$ and $i\leqslant k$ it is also contained in the strict transforms $\Delta^{i-1}$ and $S^{i-1}$. By the assumption of Theorem 0.1 the hypersurfaces $S^{+}$ and $\Delta$ meet everywhere transversally, which implies that the divisor $E_{i-1}+\Delta^{i-1}+S^{i-1}$ has normal crossings at a general point of the subvariety $B_{i-1}$, that is,
$E_{i-1}\cap\Delta^{i-1}\cap S^{i-1}$
is a non-singular subvariety of codimension 3 in $X_{i-1}$. For that reason, the more so we get $\mathop{\rm codim}(B_{i-1}\subset X_{i-1})\geqslant 3$, which completes the proof of the proposition.
Assume now that $l\leqslant k$. In that case the equality (1) can be re-written as the equality, in the left hand side of which we have the sum
$\sum^{l}_{i=1}((M+1)(m+1)-(2m+1)(\delta_{i}+e))p_{i},$
and in the right hand side we have the expression
$\sum^{k}_{i=l+1}((2m+1)\delta_{i}-M-1)p_{i}+(2m+1)\sum^{N}_{i=k+1}p_{i}\delta_{i}+(M+1).$
Since $M=2e+1$, $m\geqslant e+1$ and $\delta_{i}\geqslant 1$, the right hand side is strictly positive. However, for $i\leqslant l$ by Proposition 2.2 we have $\delta_{i}\geqslant 2$ and it is easy to check that
$(M+1)(m+1)-(2m+1)(e+2)=e-2m<0.$
Thus the left hand side is strictly negative. We obtained a contradiction that excludes the singular case for $l\leqslant k$.
2.4. Exclusion of the singular case for $l>k$. In that case the equality (1) can be re-written as the equality, in the left hand side of which we have the sum
$\sum^{k}_{i=1}((M+1)(m+1)-(2m+1)(\delta_{i}+e))p_{i}$
(if $k=0$, then it is equal to zero), and in the right hand side the expression
$\sum^{l}_{i=k+1}((2m+1)(\delta_{i}+e)-(M+1)m)p_{i}+(2m+1)\sum^{N}_{i=l+1}p_{i}\delta_{i}+(M+1).$
Thus the left hand side is non-positive (strictly negative for $k\geqslant 1$, as we saw in Subsection 2.3). Since $\delta_{i}\geqslant 1$ and the equality
$(2m+1)(e+1)-(2e+2)m=e+1$
holds, we get a contradiction again, excluding the singular case for $l>k$. This completes the proof of the claim (i) of Theorem 0.1.
2.5. Projective automorphisms, preserving the hypersurface $S$. Let us prove the claim (ii) of Theorem 0.1. We know already that $\chi_{\mathbb{P}}\in\mathop{\rm Aut}{\mathbb{P}}$ is a projective automorphism, preserving $S$. Since $\mathop{\rm Sing}S=P_{1}\cup P_{2}$, we have that $\chi_{\mathbb{P}}(P_{1})$ is either $P_{1}$ or $P_{2}$. However, by assumption $\xi_{1}(P_{1})\neq\xi_{2}(P_{2})$, so that $\chi_{\mathbb{P}}(P_{1})=P_{1}$ and $\chi_{\mathbb{P}}(P_{2})=P_{2}$. By the condition ()(*)( ∗ ) for a point of general position $p\in P_{i}$ the fibre of the exceptional divisor $\Delta_{i}$ over $p$ is not isomorphic to the fibre of $\Delta_{i}$ over any other point of the subspace $P_{i}$. Therefore
$\chi_{\mathbb{P}}|_{P_{i}}=\mathop{\rm id}\nolimits_{P_{i}}.$
Now let us consider points of general position $p_{1}\in P_{1}$ and $p_{2}\in P_{2}$ and the $[p_{1},p_{2}]$ passing through these points. The automorphism $\chi_{\mathbb{P}}$ preserves that line and its restriction $\chi_{\mathbb{P}}|_{[p_{1},p_{2}]}$ preserves the three points: $p_{1},p_{2}$ and the points
$([p_{1},p_{2}]\cap S)\backslash\{p_{1},p_{2}\}.$
Therefore, $\chi_{\mathbb{P}}|_{[p_{1},p_{2}]}=\mathop{\rm id}_{[p_{1},p_{2}]}$, so that $\chi_{\mathbb{P}}=\mathop{\rm id}_{\mathbb{P}}$, as we claimed. Q.E.D. for the part (ii).
Let us show (iii). Here we argue in a similar way with the only difference: $\chi_{\mathbb{P}}$ can swap the subspaces $P_{1},P_{2}$. If $\chi_{\mathbb{P}}(P_{i})=P_{i}$, then we can repeat the proof above word for word and conclude that $\chi_{\mathbb{P}}=\mathop{\rm id}_{\mathbb{P}}$. If $\chi_{\mathbb{P}}(P_{1})=P_{2}$, then, taking into account that the hypersurface $S$ is invariant with respect to $\tau$, we consider the composition $\tau\chi_{\mathbb{P}}\in\mathop{\rm Aut}{\mathbb{P}}$. This transformation preserves the subspaces $P_{1},P_{2}$, so that $\tau\chi_{\mathbb{P}}=\mathop{\rm id}_{\mathbb{P}}$ as above and $\chi_{\mathbb{P}}=\tau$. Q.E.D. for our theorem.